A chemical engineer wanted to etch glass with an acidic solution. The solution that he had had a 25% acid content. This is too weak. He wanted to raise the content to 40%. He added 800g of acid to the solution and raised the content to 40% What was the size of the original solution?
Answer by ankor@dixie-net.com(22740)

A chemical engineer wanted to etch glass with an acidic solution.
The solution that he had had a 25% acid content.
This is too weak. He wanted to raise the content to 40%.
He added 800g of acid to the solution and raised the content to 40%
What was the size of the original solution?
:
Let x = amt of 25% solutions
:
.25x + 800 = .40(x+800)
.25x + 800 = .40x + 320
800 - 320 = .40x - .25x
480 = .15x
x = 480/.15
x = 3200 grams of the original 25% solution